OpenStudy (anonymous):

What are the first four terms in the expansion of (2+x)^(1/2)? In what interval is the expansion valid?

OpenStudy (anonymous):

I don't understand why the interval is -2<x<2. Could someone explain that?

OpenStudy (akashdeepdeb):

(2+x)^1/2 means \[\sqrt (2+x)\] If anything is square rooted then it must be greater then or equal to 0 because if it is less than that it would be complex! SO \[2+x \ge 0\]

OpenStudy (anonymous):

Then x>=-2 why -2<x<2?

OpenStudy (akashdeepdeb):

You are right! :D The answer key is mistaken!

OpenStudy (anonymous):

Oh... Thanks!!!

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

\[\left( 2+x \right)^{\frac{ 1 }{ 2 }}=2^{\frac{ 1 }{ 2 }}\left( 1+\frac{ x }{ 2 } \right)^{\frac{ 1 }{ 2 } }\] The expansion is valid if \[\left| \frac{ x }{ 2} \right|<1, or-1<\frac{ x }{2 }<1, or -2<x <2\]

OpenStudy (anonymous):

Thanks, surjithayer.

OpenStudy (anonymous):

if you want i can expand.

OpenStudy (anonymous):

i got sqrt(2) (1+x/4-x^2/8+3x^3/16-...) the answer given by my book is different... sqrt(2) (1+x/4-x^2/32+x^3/128-...)

OpenStudy (akashdeepdeb):

@Surjith Can you expand?

OpenStudy (anonymous):

I've figured that out. But I couldn't solve another problem... Could you help? Find the coefficient of x^3 in x^(-1/2)/[2+sqrt(x)]^3 x^(-1/2)/[2+sqrt(x)]^3 = x^(-1/2) * 2^(-3) * (1+sqrt(x)/2)^(-3) (1+sqrt(x)/2)^(-3) = (1-3x^(1/2)/2+3x^(3/2)/2-5x^(5/2)/4+15x^(7/2)/16+...) (Is this right?) Coefficient of x^3: 15/16 * 2^(-3) = 15/128 The answer should be -9/256. Thanks a lot!!!

OpenStudy (anonymous):

\[\left( 1+x \right)^{n}=1+nx+\frac{ n \left( n-1 \right) }{ 2! }x ^{2}+\frac{ n \left( n-1 \right)\left( n-2 \right) }{3! }x ^{3 }+...\]

OpenStudy (anonymous):

\[n =\frac{ 1 }{ 2 }, 2!=2*1=2,3!=3*2*1=6,x=\frac{ x }{2 }\] plug the values and multiply by sqrt{2} to get the solution.

OpenStudy (anonymous):

if still problem persists i will solve

OpenStudy (anonymous):

solved. thanks :)